RC meeting and Run Plan, Wednesday November 17th

From SBS wiki
Jump to: navigation, search


RC Meeting Information

Run Coordinator: Brian Quinn

Location: Counting House Conference Room

Time: 3:15PM

Remote connection link: https://bluejeans.com/256205934

Toll-Free Number (U.S.& Canada): 1-888-240-2560

Bluejeans Meeting ID: 256 205 934

Upcoming Shifts Which Need Filled

A total of 11 shifts remain until the winter break shutdown half of which are this upcoming weekend.

  1. Sat Dec 11th, OWL, SL
  2. Sat Dec 11th, DAY, SL
  3. Sat Dec 11th, DAY, TO
  4. Sat Dec 11th, SWING, TO
  5. Sun Dec 12th, OWL, SL
  6. Sun Dec 12th, DAY, TO
  7. Wed Dec 15th, OWL, SL
  8. Fri Dec 17th, OWL, TO
  9. Sat Dec 18th, OWL, TO
  10. Sat Dec 18th, DAY, TO
  11. Sun Dec 19th, DAY, TO

Accelerator and Hall Status

  • Possible 4-hour optics studies 7 AM Tuesday.
  • 1/8" Al sheet placed outside BB vacuum window on target chamber
  • Now running with higher currents (11 uA for LD2, 15 uA for LH2)
  • Should we increase BB trig threshold to 475 mV or 480 mV (from 470)? (DT ~ 8% for LD2)


To Be Done During Extended Down Time

  • (BBCal) May potentially need more cosmics data.
  • (HCal) Would like to take special data runs ????
  • Removal of INFN J2 from Hall. (Jessie)
  • Other??

Run Plan

  • Finished Cycle 10 today (with dummy target)
  • Now taking LD2 data for Cycle 11
  • Continue following cycle pattern and ticking off runs-completed on paper copy at SL desk
  • Tomorrow? SBS-off running. 2hr LH2, 3hr LD2 (Need to change BB Cal HV settings)


Opportunistic Access Requests (Running List)

  • Change rpi & SD card in HCal's upper HV mainframe (Sebastian & Ellen)
  • Remove HV? card for hodoscope (Ralph)
  • GRINCH (change pulser rate?) (Todd/Maria)

Resolved Issues

  • "Low efficiency" of last plane of tracker
  • Low DAQ rate alarm for hung MPD (End, Reset(twice?), configure, download)
  • GRINCH LED pulser now prescaled (PS5=4) to give ~10 Hz

Unresolved Issues

  • Hodoscope HV alarms
  • Occasional timing shifts of ~70 ns


Subsystem Updates

  1. HCal
    1. ?
  2. BBCal
    1. ?
  3. Timing Hodoscope
    1. ?
  4. GRINCH
    1. ?
  5. GEMS
    1. ?
  6. Beamline
    1. ?
  7. Replay analysis
    1. ?